OpenStudy (xochital):

Which factor can you cancel from the numerator and denominator to simplify 12 minus 6 times x all over 5 times x minus 10.?

OpenStudy (ankit042):

you mean 12-6x/5x-10 ?

OpenStudy (xochital):

yes :)

OpenStudy (xochital):

could you explain to me how to solve it?

OpenStudy (ankit042):

in the numerator and denominator can you find some common factors

OpenStudy (skullpatrol):

$$\frac{12-6x}{5x-10}$$

OpenStudy (skullpatrol):

$$\frac{6(2-x)}{5(x-2)}$$

OpenStudy (xochital):

so i would factor out \[\frac{ 6 }{ 5 }\] correct?

OpenStudy (skullpatrol):

Yes, but notice also that (2-x)=-(x-2).

OpenStudy (skullpatrol):

$$\frac{6(2-x)}{5(-1)(-x+2)}$$